In the product B2×7B=6396, B is a digit. The value B =
a. 8
b. 7
c. 6
d. 5
e. None of these
: In the product B2×7B=6396, B is a digit. The value B =
a. 8
b. 7
c. 6
d. 5
e. None of these

Answer by jim_thompson5910(8311)

If we multiply B2*7B, our first step would be to multiply the last digit in 7B and the last digit in B2 to get 2*B. Since B is a single digit, we know that 2*B is going to be at most 18 (since 9 is the largest that B can be and 2*9=18). So if we subtract 10 off 2*B the largest that 2*B-10 is now 8 (which means that we now can say that 2*B-10 is equal to a single digit).


So this means that 2B-10=6



2B=6+10 Add 10 to both sides.


2B=16 Combine like terms on the right side.


B=(16)/(2) Divide both sides by 2 to isolate B.


B=8 Reduce.


----------------------------------------------------------------------

Answer:

So the answer is B=8